OpenStudy (anonymous):

whAT IS THE INVERSE OF -1/4

OpenStudy (anonymous):

4/-1

myininaya (myininaya):

additive or multiplicative?

myininaya (myininaya):

additive solve the following for a a+(-1/4)=0 multiplicative solve the following for b b(-1/4)=1

OpenStudy (anonymous):

umm i have no idea but the problem is y=-1/4x+3 (2,6) and im suppose to find the perpindicular line

myininaya (myininaya):

you need to find the opp reciprocal of the slope of that line to find the slope of the perpendicular line

OpenStudy (anonymous):

so how would i find the opposite reciprocal would that just be like 4/1

myininaya (myininaya):

yep

myininaya (myininaya):

so we have y=4x+c but you need to find the y-intercept c

myininaya (myininaya):

use the point (2,6) to find c

OpenStudy (solomonzelman):

and lastly, when you say \(\large\color{black}{ f(x)=-\frac{1}{4} }\) being that the slope is zero, there is no \(\large\color{black}{ f^{-1}(x) }\).

OpenStudy (solomonzelman):

just listing all of the definitons of "inverse".

OpenStudy (anonymous):

so the opp recprical is 4/1 and then can i simplify this number or not

myininaya (myininaya):

4/1 is 4

myininaya (myininaya):

n/1 is n

OpenStudy (anonymous):

all right i think ive got this but let me check

OpenStudy (anonymous):

so the final answer would be y=4x+-2

myininaya (myininaya):

(2,6) 6=4(2)+c 6-8=c -2=c yep y=4x-2

OpenStudy (anonymous):

yea thanks i finally understand this thank you so much

OpenStudy (jhannybean):

\[m_{\perp} = -\frac{1}{m}\]